# Assessing and Mitigating Medical Knowledge Drift and Conflicts in Large Language Models

### v2

@@
Weiyi Wu1, Xinwen Xu2, Chongyang Gao3, Xingjian Diao4, Siting Li1, Lucas A. Salas5, Jiang Gui1,4+ {affiliations}
Department of Biomedical Data Science, Dartmouth College, Hanover, NH
Medical Practice Evaluation Center, Massachusetts General Hospital, Boston, MA
Department of Computer Science, Northwestern University, Evanston, IL
Department of Computer Science, Dartmouth College, Hanover, NH
Department of Epidemiology, Dartmouth College, Hanover, NH
$\boldsymbol{+}$ Lead author
1 Abstract
Large Language Models (LLMs) have great potential in the field of health care, yet they face great challenges in adapting to rapidly evolving medical knowledge. This can lead to outdated or contradictory treatment suggestions. This study investigated how LLMs respond to evolving clinical guidelines, focusing on concept drift and internal inconsistencies. We developed the DriftMedQA benchmark to simulate guideline evolution and assessed the temporal reliability of various LLMs. Our evaluation of seven state-of-the-art models across 4,290 scenarios demonstrated difficulties in rejecting outdated recommendations and frequently endorsing conflicting guidance. Additionally, we explored two mitigation strategies: Retrieval-Augmented Generation and preference fine-tuning via Direct Preference Optimization. While each method improved model performance, their combination led to the most consistent and reliable results. These findings underscore the need to improve LLM robustness to temporal shifts to ensure more dependable applications in clinical practice. The dataset is available at https://huggingface.co/datasets/RDBH/DriftMed .
@@ 2 Introduction
Healthcare delivery is increasingly strained by the expansion of biomedical knowledge, driven by advancements in research technologies, molecular investigations, drug development, and clinical trials [1, 2, 3]. This proliferation challenges healthcare systems and practitioners, as clinicians face difficulty in staying current, particularly as their formal training may become outdated. Standard practices taught during medical education can soon become obsolete or counterproductive [4, 5], exacerbating the gap between medical education and current best practices. Empirical studies show that clinical guidelines often require reassessment within a few years [6], underscoring the need for innovative methods to support timely and reliable clinical decision-making.
Large language models (LLMs) have emerged as prospective tools for addressing this information proliferation. Recent LLMs, such as GPT‑4o [7] and Llama-3.3 [8], demonstrate physician‑level proficiency in standardized medical exams and can analyze extensive medical literature [9, 10, 11]. Healthcare systems have begun integrating these tools into clinical workflows to facilitate documentation, patient communication, and diagnostic reasoning [12]. However, despite their potential, LLMs present notable limitations. Recent studies on medical LLMs have identified biases in clinical recommendations associated with social determinants of health (SDoH) [13, 14, 15] and cognitive factors [16]. Despite growing attention to SDoH and cognitive factors in LLMs, the temporal factor associated with evolving medical knowledge remains underexplored.
The temporal progression of medical knowledge presents a fundamental challenge for LLMs in healthcare, referred to as concept drift, where established medical guidance changes over time due to the evolution of care standards. Diagnostic criteria, treatment guidelines, and management methods are regularly updated as new evidence emerges. For example, diagnostic criteria for insulin overbasalization (titration of basal insulin beyond an appropriate dose) have shifted from fixed dosage thresholds to contextual markers, including glucose trends and hypoglycemia risk [17]. Similarly, treatment guidelines for HIV/HCV co-infection have evolved from ribavirin-based regimens to newer, safer, and more effective alternatives [18].
Concept drift affects LLMs in two critical dimensions: external concept drift post-deployment and internal knowledge conflict during training.
External Medical Concept Drift: Following deployment, concept drift presents a persistent challenge for LLMs as the gap between static parametric knowledge and the changing medical reality continuously widens. As medical advancements surpass a model’s fixed understanding, concept drift gradually erodes model trustworthiness. Static models developed on the fixed dataset cannot adapt to emerging health threats or incorporate new findings after training, leading to growing misalignment with current clinical practice [19]. This divergence becomes obvious during public health emergencies, such as the COVID-19 pandemic, when the infection control protocols underwent frequent and major revisions, and sticking to outdated recommendations posed increased risks [20, 21].
Internal Medical Knowledge Conflict: During training, LLMs face challenges processing heterogeneous and inconsistently curated data from different periods without the ability to distinguish temporal context [22, 23]. Research demonstrates that LLMs have intrinsic difficulties with training corpora: temporal misalignment, embedded disinformation, and inherent biases [24]. This risk is particularly acute in medicine, where evolving standards are common. LLMs may treat outdated and contemporary recommendations without mechanisms to resolve temporal inconsistencies as equally valid, potentially leading to inappropriate or harmful suggestions. This challenge is further compounded when clinical guidelines undergo reversals over time [25, 26]. For example, the NICE-SUGAR study demonstrated that intensive glucose control in critically ill patients, once thought beneficial based on earlier evidence, actually increased mortality compared to conventional control [25]. Such guideline reversals underscore the urgent need for LLMs not only to incorporate updated evidence but also to unlearn outdated standards that have been shown to be ineffective or dangerous.
Figure 1: Overview of DriftMedQA benchmark construction and prompt example. (Left) Up-to-date clinical guidelines are paired with manually constructed pseudo-outdated counterparts. Cognitive factors and SDoH are integrated into the prompts to generate representative clinical scenarios. (Right) Example advice pairs showing raw guideline content used in the scenario construction. (Bottom) Example of a final model evaluation prompt containing a contextual narrative with embedded self-diagnosis bias.
To systematically assess how internal medical knowledge conflicts and external medical concept drift affect LLM performance and reliability, we developed DriftMedQA, a benchmark designed to simulate the temporal evolution of clinical knowledge. DriftMedQA pairs up-to-date clinical recommendations with manually constructed pseudo-outdated counterparts, enabling direct evaluation of model alignment with evolving standards. To better reflect real-world clinical decision-making complexity, we employed LLMs to generate diverse clinical scenarios based on medical advice, optionally incorporating cognitive factors or SDoH. These enriched contexts simulate realistic patient presentations. A schematic overview of the benchmark construction process and example prompt is provided in Figure 1.
Using this enriched framework, we evaluated seven cutting-edge LLMs across a diverse set of 4,290 annotated clinical scenarios. We provided every LLM with the same prompt to control for variability and ensure consistency for direct comparison, enabling us to quantify how temporal knowledge drift and internal knowledge conflicts affect their reliability. Our baseline findings demonstrate that LLMs consistently struggle to detect when medical knowledge has changed, often failing to embrace up-to-date guidance or to flag flawed, outdated recommendations. Models also exhibited systematic vulnerability to internal conflicts, particularly when encountering subtle but clinically critical changes, where both the old and new recommendations could be misinterpreted as equally current.
To begin addressing these challenges, we explored two complementary mitigation strategies. Retrieval-Augmented Generation (RAG) was employed to incorporate up-to-date external knowledge at inference time, and preference fine-tuning (using Direct Preference Optimization, DPO) was used for parameter-efficient fine-tuning to prioritize correct contemporary advice over outdated standards. While RAG and DPO individually improved model performance, their combined application yielded the greatest benefits, significantly enhancing alignment with current guidelines and reducing internal contradictions. Nevertheless, even with these mitigations, LLMs remained susceptible to reasoning errors in complex clinical environments. Our findings highlight the critical need for continuous updating mechanisms, robust temporal evaluation frameworks, and hybrid strategies that combine external retrieval with internal adaptation to support LLMs’ safe and reliable use in dynamic healthcare settings.
Our work contributes to the growing body of evidence on how knowledge conflicts arise in LLMs and explores preliminary methods for dynamically updating models in response to evolving evidence. Although more robust mechanisms for resolving knowledge drift and reconciling conflicting information are still needed, these findings offer a crucial step toward safer and more effective LLM deployment in healthcare. This work underscores the importance of ongoing efforts to develop dynamic, temporally aware LLM systems by illuminating the pitfalls of static knowledge and knowledge conflicts in an ever-changing field.
3 Results
3.1 Model evaluation
We first assessed the baseline performance of LLMs on the DriftMedQA benchmark. Seven state-of-the-art LLMs were evaluated across 4,290 synthetic clinical scenarios designed to test temporal reliability. To capture diverse patterns of guideline evolution, in the initial evaluation phase, model performance was analyzed across five clinically meaningful update dimensions: Clinical Context, Diagnostic Thresholds, Implementation Approach, Recommendation Intensity, and Treatment Modality. Performance was assessed using three complementary metrics: endorsement of current clinical advice (ECDAadh), rejection of outdated advice (ECDArej), and overall alignment score (ECDAall); see the Methods section for details.
@@ 3.1 Model evaluation
All assessed models exhibited varying performance across the five types of clinical recommendation updates (Figure 2). GPT-4o and Qwen2.5-72B demonstrated the highest endorsement scores (ECDAadh), with sample-weighted averages of 0.90 and 0.92, respectively. These scores were significantly higher than the third-best performing model, Qwen2.5-7B (both $p<0.0001$). However, both models exhibited substantial declines when assessed on their ability to reject pseudo-outdated recommendations (ECDArej), with scores of 0.395 for GPT-4o and 0.278 for Qwen2.5-72B. Conversely, Ministral-8B achieved the highest rejection performance (ECDArej $=0.80$), following by gemma-2-27B (ECDArej $=0.68$) and Llama-3-8B (ECDArej $=0.63$), outperforming models with stronger endorsement capabilities (Figure 2(b)).
When considering overall alignment across both current and outdated scenarios, GPT-4o achieved the highest score (ECDAall $=0.65$), as shown in Figure 2(c). This performance was significantly higher than that of the second-best model, Llama-3.3-70B (ECDAall $=0.61$, $p=0.00033$), and the third-best model, Qwen2.5-72B (ECDAall $=0.60$, $p=0.0006$).
@@
As shown in Figure 3, all evaluated models exhibited measurable internal conflicts, with considerable variability across models and types of guideline updates. Our analysis revealed that more capable or larger-scale models did not consistently exhibit lower internal knowledge conflict ratios (IKCRs; see the Methods section for details). For instance, the 72B parameter version of Qwen2.5 demonstrated higher IKCRs than its 7B counterpart across most evaluated categories. Similarly, Llama-3.3-70B did not show lower conflict ratios compared to Llama-3-8B. Among all models evaluated, the Mistral-8B model achieved the lowest overall IKCR, with a weighted average score of 0.34 across all scenario types, followed by Gemma-2-27B at 0.39.
All evaluated models exhibited knowledge conflicts across all five modification categories. The highest average IKCRs were observed for changes under the groups Implementation Approach (average IKCR $=0.55$) and Treatment Modality (average IKCR $=0.53$). While our baseline evaluation distinguished performance across five guideline change categories, the mitigation analysis focuses on overall alignment and conflict rates to emphasize aggregate improvements.
3.2 Mitigation strategies
We evaluated two complementary strategies to improve temporal reliability and internal consistency in LLM outputs: Retrieval Augmented Generation (RAG) and Preference fine-tuning via Direct Preference Optimization (DPO), as illustrated in Figure4. Each method was tested individually and in combination across different LLMs (see the Methods section for details), with model performance assessed on ECDA and IKCR metrics as summarized in Figure 5. While not exhaustive, these initial strategies helped illuminate where and how to improve temporal alignment.

### v3

@@
Abstract
Large Language Models (LLMs) offer transformative potential across diverse fields, yet their safe and effective deployment is hindered by inherent knowledge conflicts—stemming from temporal evolution, divergent sources, and contradictory guidelines. This challenge is particularly acute in medicine, an interdisciplinary frontier for NLP. Rapid medical concept drift can lead LLMs to provide incorrect or outdated advice, impacting their utility and the broader societal benefits of NLP advances. This study introduces ConflictMedQA, a benchmark designed to systematically evaluate how LLMs manage varied knowledge conflicts in clinical guidelines. Our assessment of seven state-of-the-art models across 4,290 scenarios reveals significant difficulties in rejecting incorrect recommendations and frequent endorsement of conflicting advice, highlighting an important gap for NLP systems intended for real-world impact. We explore two fundamental mitigation approaches: retrieval-augmented generation and preference fine-tuning via direct preference optimization. While each offers improvements, their synergistic combination yields the best results. These findings emphasize the need for LLMs to discern subtle but critical guideline conflicts. This is a crucial step in advancing NLP’s capabilities and ensuring its dependable application in critical societal domains. The proposed dataset is available at https://huggingface.co/datasets/RDBH/DriftMed.
@@ Abstract
Weiyi Wu1, Xinwen Xu2, Chongyang Gao3, Xingjian Diao1, Siting Li1, Lucas A. Salas1, Jiang Gui1 1Dartmouth College, 2Massachusetts General Hospital, 3Northwestern University weiyi.wu.gr@dartmouth.edu
Figure 1: Overview of ConflictMedQA benchmark construction and prompt example. (Left) Up-to-date clinical guidelines are paired with manually constructed pseudo-outdated counterparts. Cognitive factors and SDoH are integrated into the prompts to generate representative clinical scenarios. (Right) Example advice pairs showing raw guideline content used in the scenario construction. (Bottom) Example of a final model evaluation prompt containing a contextual narrative with embedded self-diagnosis bias.
@@ 1 Introduction
The rapid expansion of biomedical knowledge, driven by swift research and medical advancements, increasingly strains healthcare delivery Densen (2011); Chopra et al. (2023); Singh et al. (2023). Clinicians struggle to stay current as standard practices can quickly become obsolete Lajoie and Gube (2018); Halalau et al. (2021), with clinical guidelines—the formal standards of medical knowledge—often needing reassessment within years Shekelle et al. (2001). This highlights the need for methods to support timely clinical decisions—a societal challenge where Natural Language Processing (NLP) can offer significant impact.
Large Language Models (LLMs) are promising tools to navigate this information, showing strong clinical text comprehension and reasoning Tu et al. (2025); Singhal et al. (2025); Liévin et al. (2024); Singhal et al. (2023). While healthcare explores their integration Thirunavukarasu et al. (2023); Glicksberg et al. (2024), their transformative potential hinges on rigorously understanding limitations beyond exam accuracies. Research has widely explored clinical biases in LLMs Zack et al. (2024); Schmidgall et al. (2024). Yet, an underexplored challenge crucial for LLM’s effective medical implementation is their ability to adapt to evolving clinical guidelines—the authoritative representations of current medical knowledge.
This guideline evolution creates two challenges. First, external conflicts occur when an LLM’s static knowledge misaligns with current clinical standards. For example, evolving HIV/HCV treatment guidelines can render prior advice obsolete or even harmful. Second, internal knowledge conflicts arise when LLMs assimilate contradictory guidelines from diverse training data Xie et al. (2024); Chen et al. (2022). The NICE-SUGAR study on glucose control Investigators (2009); Cagnacci and Venier (2019) exemplifies this challenge, where intensive glucose management, once recommended in guidelines, was later found to increase mortality. Such guideline reversals erode trust and impede NLP’s impact when LLMs provide contradictory advice Abdool Karim and Devnarain (2022); Jean and Hsueh (2020).
Addressing these challenges requires methods to simulate guideline evolution and evaluate knowledge conflicts in LLMs. Current medical benchmarks predominantly focus on static knowledge and well-established facts, neglecting how knowledge evolves over time. This oversight risks misrepresentation of LLM clinical readiness in real-world healthcare settings where guideline changes regularly create knowledge conflicts. We developed ConflictMedQA (Fig. 1), a benchmark that simulates guideline evolution to assess how LLMs manage conflicts between previous and current medical knowledge standards. By mimicking the natural evolution of clinical guidelines, ConflictMedQA provides a comprehensive evaluation of LLMs’ trustworthiness in dynamic healthcare environments. This work’s contributions are:
  • We introduce ConflictMedQA, a benchmark assessing LLMs’ handling of resulting knowledge conflicts in healthcare.
  • Our empirical analysis reveals LLM limitations in reconciling conflicting medical knowledge, highlighting gaps in clinical readiness.
  • We propose a framework combining two strategies that provides a promising way for improving LLM adaptation to evolving medical knowledge.
2 Related Works
2.1 LLMs in Healthcare
LLMs show remarkable capabilities, with healthcare a prominent application area. Models like GPT-4o Achiam et al. (2023) and Llama 2 Touvron et al. (2023) show physician-level proficiency on medical exams and can synthesize medical literature Singhal et al. (2025); Liévin et al. (2024); Singhal et al. (2023). This spurs interest in their clinical integration for documentation, patient communication, and diagnostic aid Thirunavukarasu et al. (2023). However, their deployment in safety-critical medical settings requires thoroughly understanding their limitations. Social determinants of health (SDoH) and cognitive factors have been shown to influence both real-world clinical decision making Ma et al. (2025); Hammond et al. (2021) and LLM-generated recommendations Zack et al. (2024); Schmidgall et al. (2024); Liu et al. (2024). Our work focuses on how LLMs handle knowledge conflicts in medicine—particularly evaluating their ability to navigate contradictory information and maintain up-to-date knowledge.
@@ 2.1 LLMs in Healthcare
2.2 Knowledge Conflicts and Concept Drift
Training LLMs on vast, diverse, and temporally varied datasets containing contradictory information can cause internal knowledge conflicts Chen et al. (2022); Xie et al. (2024), where models hold mutually exclusive information. The conflict issue is exacerbated as models tend to memorize their training data rather than learning to generalize or resolve contradictions Yuan et al. (2025). Xu et al. Xu et al. (2024) explored identifying and resolving such conflicts in general LLMs, stressing factual consistency. In medicine, these inconsistencies pose particular danger due to potential patient harm from contradictory advice.
Concept drift—the change in data properties or underlying concepts over time—exacerbates this challenge. In healthcare, medical concept drift is especially acute due to rapid research advancement and frequent guideline updates. Public health crises like COVID-19 highlighted this vulnerability, with information evolving daily Abdool Karim and Devnarain (2022); Jean and Hsueh (2020). Guideline reversals, where previously recommended practices are later found harmful, create significant knowledge conflicts when LLMs ingest both old and new recommendations without proper prioritization Investigators (2009); Cagnacci and Venier (2019).
These problems are complicated by shifts in diagnostic criteria that move beyond simple thresholds to more nuanced, contextual markers reflecting deeper pathophysiological understanding Committee (2025), and treatment protocols that evolve toward safer, more effective regimens. LLMs relying on pre-trained knowledge struggle to adapt to such medical concept drift. Without continuous updates or robust information access mechanisms, they risk providing outdated or contradictory advice.
@@ 2.2 Knowledge Conflicts and Concept Drift
3 Evaluations
3.1 Benchmark Construction
We developed ConflictMedQA, a dataset of 195 clinical recommendation pairs covering infectious (n = 66) and chronic diseases (n = 129). Each pair includes current recommendations alongside manually created, mutually exclusive, pseudo-outdated versions. We derived these pseudo-outdated recommendations using five strategies reflecting common patterns of knowledge evolution in clinical guideline updates:
  • Clinical Context (N=22, 11.3%): Revisions to the specific patient populations or clinical circumstances to which a recommendation applies (e.g., narrowing or broadening age ranges).
  • Diagnostic & Threshold (N=42, 21.5%): Modifications to specific numerical criteria or classifications used in diagnosis or risk stratification (e.g., changing diagnostic thresholds).
  • Implementation Approach (N=32, 16.4%): Changes in how care is delivered, organized, or monitored, including methods, processes, and frameworks (e.g., shifting from one mode of care delivery or monitoring to another).
  • Recommendation Intensity (N=53, 27.2%): Changes in the strength or certainty of a recommendation while the core action remains the same (e.g., shifting from permissive to directive language).
  • Treatment Modality (N=46, 24.6%): Changes in the specific medical interventions recommended (e.g., replacing an older drug class with a newer one).
To further evaluate LLM performance under clinically relevant and cognitively diverse conditions, we transformed each medical recommendation into a richly contextualized, scenario-based question-answer (QA) pair. This design was motivated by prior work highlighting the impact of cognitive biases and SDoH on LLM clinical reasoning Schmidgall et al. (2024); Zack et al. (2024). Each scenario was conditioned on one of ten cognitive or social factors commonly encountered in medical decision-making, with an additional neutral “No Factor" setting in which no cognitive factor or SDoH was introduced. The selected factors—self-diagnosis, recency, confirmation, frequency, status quo, cultural, socioeconomic, racial or ethnic, geographical, and false consensus — capture realistic variations in reasoning without introducing factual distortion or adversarial intent.
We used Qwen2.5-72B Yang et al. (2024) to generate these scenarios by systematically combining each medical recommendation with its corresponding factor. This pipeline produced a total of 4,2904{,}290 scenario-based QA pairs $(11~\text{factors}\times 195~\text{recommendation}\times 2~)$, evenly split between current and wrong recommendations.
3.2 Models & Evaluation Metrics
We evaluated seven LLMs spanning a range of model sizes and architectures: GPT-4o Achiam et al. (2023), Llama-3-8B-Instruct and Llama-3.3-70B-Instruct Touvron et al. (2023), Qwen2.5-7B-Instruct and Qwen2.5-72B-Instruct Yang et al. (2024), Gemma-2-27B-it Team et al. (2024), and Ministral-Instruct Jiang et al. (2024). Detailed descriptions are provided in the appendix.
We evaluated LLMs’ clinical reliability through two complementary dimensions: one quantifying conflicts with external evolving medical guidelines and the other detecting internal knowledge inconsistencies.
External Knowledge Conflicts: To quantify model alignment with external evolving medical guidelines, we assess model performance across temporally distinct medical scenarios. This is measured by a set of metrics we term External Concept Drift Alignment (ECDA). Let $\mathcal{D}_{U}$ denote the set of up-to-date scenarios (n=2,145n=2{,}145) where endorsement is the correct action, and $\mathcal{D}_{O}$ represent outdated scenarios (n=2,145n=2{,}145) where rejection is appropriate. For each scenario $s_{i,c,t}$ — representing concept $i$, change type $c$, and temporal status $t\in\{u,o\}$. Let $\hat{y}_{i,c,t}\in\{0,1\}$ denote the model’s binary prediction ($1=\text{endorse, }0=\text{reject}$) and $y_{i,c,t}$ the ground truth (1 if $t=u$, 0 if $t=o$). We define alignment metrics as follows:
$\text{ECDA}_{\text{adh}}$ (Eq. 1) measures the model’s ability to correctly endorse current medical guidelines ($y_{i,c,u}=1$), while $\text{ECDA}_{\text{rej}}$ (Eq. 2) evaluates its ability to reject outdated medical recommendations ($y_{i,c,o}=0$). Their average $\text{ECDA}_{\text{all}}$ (Eq. 3) provides a balanced assessment of external conflicts with the current guidelines.
Internal Knowledge Conflicts: To detect internal knowledge inconsistencies, we evaluated whether models simultaneously endorsed conflicting recommendations using the Internal Knowledge Conflict Ratio (IKCR). Our evaluation scenarios present paired current ($s_{i,c,u}$) and outdated ($s_{i,c,o}$) versions for each core clinical concept $i$ and change $c$. Let $\hat{y}_{i,c,u}$ and $\hat{y}_{i,c,o}$ be the model’s binary predictions ($1=\text{endorse}$). We define the set of active pairs, $\mathcal{A}$, as those where the model endorses at least one version ($\mathcal{A}=\{(i,c)\mid\hat{y}_{i,c,u}=1\lor\hat{y}_{i,c,o}=1\}$). An internal contradiction, or knowledge conflict, occurs for an active pair $(i,c)\in\mathcal{A}$ when the model simultaneously endorses both mutually exclusive recommendations ($\hat{y}_{i,c,u}=1\land\hat{y}_{i,c,o}=1$). The IKCR quantifies the frequency of such contradictions:
4 Mitigating Strategies
We explored three strategies to address this challenge: non-parametric knowledge update, parametric knowledge adaptation, and hybrid knowledge augmentation. Non-parametric update was applied to all evaluated LLMs. Due to limited training resources and lack of access to proprietary model weights, parametric and hybrid knowledge update strategies were evaluated only on Qwen2.5-7B, Ministral-8B, and Llama-3-8B.
4.1 Non-Parametric Knowledge Update
This strategy supplements the model with external information during inference without modifying its internal parameters. Specifically, we employed Retrieval-Augmented Generation (RAG) Lewis et al. (2020), using a knowledge base of 195 up-to-date clinical advice.
For each clinical query scenario $s$, we encode the query using Sentence-BERT encoders Reimers and Gurevych (2019); Wang et al. (2020) and retrieve the top-$k$ most relevant guideline snippets ($d_{i}$) from our knowledge base ($\mathcal{KB}$) based on cosine similarity, then augment the input prompt with these documents before generating the response:
$\hat{y}_{s}$ $=\text{LLM}\!\bigl{(}s\oplus D_{k};\theta_{\text{base}}\bigr{)}.$ (6)
where $E_{q}$ and $E_{d}$ are query and document encoders, $\cos$ denotes cosine similarity, $k=2$ in our experiments, $\oplus$ represents prompt concatenation, and $\theta_{\text{base}}$ denotes the unchanged base model parameters. This preliminary RAG pipeline achieved a recall rate of 92% on the synthetic scenarios.
This non-parametric strategy delivers clear clinical benefits: it decouples the model from its knowledge source, allows guideline updates to be incorporated instantly without retraining, and retains explicit citations to authoritative documents. Those advantages, however, come with costs. The knowledge base demands continual curation and governance; each inference step triggers a retrieval call, adding latency and operational complexity; system performance depends on the coverage and freshness of external sources; and retrieval errors can introduce hallucinations or amplify existing biases.